# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v3

@@ III Signal Detection
This section discusses the application of deep unfolding techniques for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
@@ III-A Massive MIMO
The projected gradient descent (PGD) method is a popular optimization-driven approach for MIMO detection. In each iteration, PGD performs gradient descent to minimize detection error, then projects the solution onto a feasible, valid constellation set [48]. One of the first papers to use deep unfolding in detection is the DetNet architecture [49]. To achieve ML optimization, the PGD algorithm is unfolded where the gradient’s step size is learned across layers. The authors in [49] claimed that the computational complexity is reduced compared to the SD. DetNet nearly matches SD’s performance (within 0.5 order of BER at 18 dB). The computational complexity is linear ($\mathcal{O}(L.K^{2})$), where $L$ and $K$ are the number of layers and constellation size, respectively. On the other hand, the SD exhibits ($\mathcal{O}(2^{K})$) for worst-case complexity. However, the proposed network is not flexible to changes in the size of the constellation or the number of users. To address MIMO detection under imperfect CSI, He et al. introduced model-driven DL networks, orthogonal approximate message passing (OAMP) network, OAMP-Net [10], and its enhanced variant OAMP-Net2 [11]. OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions. The OAMP algorithm is unfolded, and the linear MMSE (LMMSE) channel estimator is used to design the receiver in [11]. Learnable parameters, such as the step-size for updating the prior mean and variance in the MMSE estimator, are trained using SGD. These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [49].
@@
As the number of transmit antennas exceeds the number of receive antennas, low complexity detectors are required for such overloaded massive MIMO systems. In [50], the trainable PGD (TPGD) algorithm was introduced for such systems. TPGD uses a DL-aided iterative PGD method with trainable parameters, such as step-size, which are optimized via backpropagation and SGD. TPGD achieves lower computational cost with fewer trainable parameters compared to DetNet [49], OAMP-Net2 [10]. Yun et al. in [51] enhanced the unfolded PGD design by adding momentum components to the iterative hidden layers. Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [52]. This modification improves convergence speed and overall performance, outperforming DetNet [49] and OAMP-Net2 [11]. Furthermore, a trainable detector based on the Hubbard–Stratonovich transformation (to avoid matrix inversion) is discussed in [53]. Compared to the TPGD detector [50] and OAMP-Net2 [11], the network in [53] achieves faster training and lower execution cost. In [54], the authors introduced the MIMO detection network (MMNet), an unfolded detector designed for ill-conditioned channels and non i.i.d. scenarios. MMNet is based on an iterative shrinkage-thresholding algorithm (ISTA), alternating between a linear detector and a nonlinear denoising step. The architectural flexibility and computational complexity are balanced, outperforming DetNet [49] and OAMP-Net [10]. The MMNet outperforms the classical LMMSE detector with a performance gain of approximately $4-6$ dB at a symbol error rate (SER) of $10^{-3}$ for different channel conditions.
@@
In [55], the CONCRETE MAP method is proposed, based on a CONtinuous relaxation of the probability mass function for disCRETE random variables, as derived in [56]. This approach solves the maximum a posteriori (MAP) detection problem using continuous optimization instead of exhaustive search, approximating the MAP solution for any probabilistic nonlinear model. By unfolding the gradient descent algorithm, the detector is able to generate soft outputs, making it suitable for integration with soft-input decoders. The network in [55] outperforms DetNet [49], OAMP-Net2 [11], and MMNet [54], in terms of BER under correlated channels. In [57], conjugate gradient descent (CGD) iterations are unfolded to design the network, LcgNet. The authors also proposed a quantized version, QLcgNet, which minimizes memory usage with negligible performance degradation. LcgNet outperforms DetNet [49] in terms of SER, computational complexity, and memory cost, making it a robust solution for massive MIMO systems. To enhance LcgNet, Olutayo et al. [58] incorporated a preconditioner during training, designing a preconditioned LcgNet. Preconditioning improves the condition number of a matrix [59], optimizing the receiver filter’s spectrum. This enhancement significantly increases the convergence speed compared to the original LcgNet in [57]. Additionally, the dynamic conjugate gradient network utilizes self-supervised learning with forward error correction codes, effectively adapting to non-stationary, time-varying channels.
@@
Physical layer signal processing involves continuous-time analog waveforms, which must be quantized for hardware implementation. Low-resolution (LR) one-bit analog-to-digital converters (ADC) are commonly used at radio frequency (RF) ends due to their high sampling rates, low cost, and low energy consumption [60]. To recover high-dimensional signals from noisy measurements, Khobahi et al. [61] developed an LR detection network (LoRDNet). This network extracts signals from noisy one-bit measurements with unknown CSI. It is based on unfolding the PGD method for symbol detection using the ML estimator. The channel effects are embedded in the training data, so LoRDNet implicitly learns a representation of the channel behavior, eliminating the need for CSI. The network outperforms classical ML estimators [62] and DL networks like DeepSIC [63], even with small datasets of about 500 samples. In [64], the authors proposed a generalized PGD algorithm. The generalized PGD algorithm executes multiple gradient descent steps before applying the projection, allowing for a better balance between convergence speed and detection performance. The authors introduced the self-correcting auto detector (SAD) by unfolding the generalized PGD algorithm. The architecture has two modules: the self-correcting module that executes multiple gradient descent steps to refine the detection estimate, and a denoising autoencoder, a learnable projection module. The numerical results demonstrate that SAD [64] achieves a better trade-off between detection accuracy and complexity compared to DetNet [49]. Although DetNet requires $15$ projection steps, SAD only performs $4$, significantly reducing detection latency and computational cost.
The channel matrix for a MIMO fading channel can be modeled as a graph. Graph models, such as Bayesian belief networks, Markov random fields, and factor graphs, combine the posterior probability of transmitted data variables and graph theory, establishing a relationship between different nodes for proper detection [3]. The performance of these methods largely depends on the underlying graph structure and convergence properties. Iterative algorithms, like message-passing detectors (MPD) with belief propagation (BP), approximate message passing (AMP), and channel hardening-exploiting message passing (CHEMP), further reduce computational complexity. Expectation propagation (EP), as an extension of BP, enables efficient detection in higher-dimensional MIMO systems and large constellations. EP approximates the posterior distribution of transmitted symbols iteratively, enabling efficient detection in polynomial time [65, 66]. Tan et al. [67] unfolded MPD iterations to update messages at symbol nodes using prior and posterior log-likelihood ratio (LLR). Learnable parameters, such as prior probabilities and correction factors (damping, scaling factor, LLR rescaling), are optimized via mini-batch SGD. The authors unfolded the existing classical MPDs: BP and CHEMP, where the correction factors are learned instead of heuristic tuning. A greedy search determines the number of layers. The unfolded networks adapt better to correlated channels and perform consistently in higher-order modulation than classical MPDs. In [68], a low-complexity AMP-graph NN (GNN) model (AMP-GNN) is proposed, which outperforms OAMP-Net2 [11] by avoiding matrix inversion. Due to the high computational complexity, message-passing algorithms are generally unsuitable for single-carrier frequency division multiple access (FDMA) systems, where a low peak-to-average power ratio is a key design objective for efficient transmission. To address this, the unfolded network in [69] simplifies generalized AMP iterations as discussed in [70]. The proposed network simplified generalized AMP iterations by approximating per-subcarrier channel statistics through statistical averaging. The unfolded network [69] maintains high performance even with channel mismatches and achieves $0.83$ dB at BER$=10^{-3}$ gain over classical generalized AMP [70]. To enhance EP-based detection, the authors in [71] proposed a network, a modified EP MIMO detector (MEPD). In the network, parameter selection is optimized using a modified moment-matching scheme and shows superior performance in high modulation cases compared to OAMP-Net2 [11]. In [72], an approximate EP inspired by [73] is unfolded, achieving better detection performance and robust convergence.
@@
TABLE III: Summary of typical deep unfolding models for signal detection and their advantages and disadvantages.
Unfolding Model Main Novelties and Performance Gains Advantages and Disadvantages
DetNet [49] First deep unfolding model for MIMO signal detection based on unrolling PGD method. In a $60\times 30$ MIMO system, DetNet achieves $3$ dB SNR gain over ZF for BER $=10^{-3}$, performs comparable to SDR and AMP, and approaches SD. Advantages: Explainable (follows PGD optimization flow), performs well when $N_{r}\gg N_{t}$. Disadvantages: Complex architecture (high computational cost), poor generalization under correlated channels.
LcgNet [57] Unfolded CGD for massive MIMO detection. Achieves $1.52$ dB SNR gain over LMMSE in $32\times 128$ MIMO at BER $=10^{-4}$. Outperforms DetNet with lower complexity and faster training. Advantages: Fast training (practical for low-end CPUs). Disadvantages: Trained on Rayleigh fading, suffers significant BER degradation when tested on a spatially correlated channel.
OAMP-Net2 [11] Lightweight unfolding of OAMP with soft input/output. Outperforms DetNet and AMP in 4$\times$4 and 8$\times$8 MIMO system, robust under imperfect CSI. Advantages: Fast training (4 parameters). Supports turbo equalization (soft I/O). Disadvantages: Underperforms in 3GPP channels (assumes unitarily-invariant matrix).
MMNet [54] Online-trained unfolded ISTA for 3GPP channels. Achieves $8$ dB SNR gain over MMSE for SER $=10^{-3}$ (and $2.5$ dB over OAMP-Net2 [11] under imperfect CSI). Advantages: Robust to correlated channels, high-order QAM, and estimation errors. Disadvantages: Online training adds latency (unsuitable for fast-varying channels).
ADMM-Net [38] Inexact ADMM-based unfolding with soft $tanh$ demodulation (supports $16/64$ QAM). Achieves SER within $1.2$ dB of SD in $30\times 60$ MIMO, outperforms DetNet with lower complexity. Advantages: Robust to channel correlation and imperfect CSI. Disadvantages: More layers (5205205–205 – 20) improve accuracy but increase computation, hence a tradeoff between performance and runtime.
LoRDNet [61] Unfolding PGD for 1-bit ADC receivers without CSI. Outperforms classical ML detectors with just 500 samples. Advantages: Works with low-resolution ADCs; data-efficient. Disadvantages: Periodic retraining required for block fading channels (reduced SE due to pilot overhead) .
@@
To avoid the computational burden incurred due to matrix inversion, the authors in [78] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO. The authors incorporated deep unfolding for components in matrix inversion, to reduce the detection complexity. The sparse weight matrix associated with the Jacobian iteration is trained across the layers of the network. Liao et al. [32] incorporated interference cancellation into DNN layers. Inspired by [79], their sequential detector trains residual error coefficients with mean square error (MSE) loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication. Compared to DetNet [49], which requires $90$ layers for convergence, their model achieves convergence with just $8$ layers, offering significant computational simplicity. The authors in [38] developed an unfolded ADMM-based network for MIMO detection. They employed an inexact ADMM update as in [37, 80] to reduce complexity while achieving near-optimal performance. Compared to DetNet [49], the proposed network [38] requires fewer trainable parameters, resulting in improved performance in terms of symbol error. Since, in [38], perfect CSI was assumed for MIMO detection, the authors in [81], proposed an ADMM-based unfolded algorithm with channel estimation errors. A set of penalty and relaxation parameters are learned layer wise. Moreover, the complexity of the proposed network is reduced as compared to OAMP-Net2 [11].
@@ III-B Modern Multicarrier Modulation
A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [82]. The current trend in wireless communication is to increase the signal bandwidth to achieve higher transmission rates. However, increasing bandwidth results in shorter symbol periods due to the inverse relationship between bandwidth and symbol period. Consequently, in a multi-path fading channel, the reduced symbol period exacerbates ISI. Multi-carrier modulation techniques like OFDM provide an effective solution to mitigate ISI by dividing a wideband signal into a set of orthogonal narrowband subcarriers [83]. MIMO-OFDM detectors face many challenges due to the complexity introduced by high dimensional signal space. A low complexity receiver for a cyclic prefix-free MIMO-OFDM system is explored in [84]. The receiver unfolds the OAMP algorithm and scalable parameters are introduced for optimizing the channel estimate matrix. The network was tested in over the air experiments and it performed reliably with low synchronization errors under imperfect CSI. Similarly, detection of coarsely quantized or LR signals in one-bit OFDM-based MIMO systems is a challenge. Hence, to address this challenge Shao et al. [85] have proposed an unfolded expectation maximization (EM) algorithm. One-bit ADCs severely quantize received signals, discarding amplitude information, and hence, ML detection becomes intractable due to nonlinear quantization effects. Hence, EM iteratively estimates unquantized symbols by computing its conditional expectation and optimizes the parameters in the maximization step. At high SNR, unfolded EM [85] performs better than ZF for one bit MIMO-OFDM detection. Furthermore, by unfolding the likelihood ascent search algorithm, Ullah et al. [86] developed an unfolded network to jointly detect and estimate the soft output in MIMO-OFDM systems. The likelihood ascent search algorithm is a neighborhood search-based detector for the MIMO system. The algorithm begins with an initial symbol estimate from linear detectors (ZF or MMSE) and iteratively refines the solution by flipping bits in the detected symbols to minimize the likelihood cost function. Near-optimal detection is achieved with the unfolded network, $1.2$ dB gap with 16-QAM as compared to SD at BER of $10^{-5}$ at polynomial complexity. Zhang et al. developed an unfolding-based framework for MIMO-OFDM detectors in [87]. The EP is unfolded with layer-specific damping factors that are trained to adjust effectively to varying channel conditions. The proposed network achieved near-optimal detection across a wide range of channels and noise levels and demonstrated robustness in over-the-air experiment tests. Underwater acoustic communication is widely used for long-range transmissions, reaching up to tens of kilometers [88, 89, 90]. Sound waves are used to transmit data through water since radio and light waves attenuate rapidly underwater. Channel estimation in underwater acoustic communication is particularly challenging due to doubly-selective channels that exhibit both severe time variation (from Doppler effects) and frequency selectivity (from long multipath delays). Hence, for OFDM systems in complex doubly-selective underwater acoustic channels, Zhao et al. in [35] proposed an unfolded equalizer (UDNet). The unfolded algorithm is based on unfolding the classical MMSE equalizer. In terms of performance, UDNet achieved better generalization and maintained lower computational complexity than DetNet [49]. Li et al. [34] proposed an unfolded EP network (EP-Net) for OTFS detection. The trainable parameters include the step-size, damping factor, and offset values for the mean and variance of the posterior conditional probability distribution. To enhance convergence and prevent the network from being trapped in local optima, the Adam optimizer with warm restarts is employed. In [91], an unfolded OAMP-based detector for OTFS is designed to handle channel estimation errors. This detector is trained using imperfect CSI and optimizes three trainable parameters: step-sizes for updating the linear estimate, nonlinear estimate, and noise variance. Numerical analysis demonstrates that at SNR of $30$ dB, the proposed detector achieves a BER of $2.4\times 10^{-4}$ while EP-Net [34] achieves BER of $1.2\times 10^{-3}$.

### v4

@@ III Signal Detection
This section discusses the applications of deep unfolding for signal detection. We segregate these signal detection techniques into various key technologies, as described in the following.
@@ III-A Massive MIMO
The projected gradient descent (PGD) method is a popular optimization-driven approach for MIMO detection. In each iteration, PGD performs gradient descent to minimize detection error, then projects the solution onto a feasible, valid constellation set [184]. One of the first papers to use deep unfolding in detection is the DetNet architecture [163]. To achieve ML optimization, the PGD algorithm is unfolded where the gradient’s step size is learned across layers. The authors in [163] claimed that the computational complexity is reduced compared to the SD. DetNet nearly matches SD’s performance (within 0.5 order of BER at 18 dB). The computational complexity is linear ($\mathcal{O}(L.K^{2})$), where $L$ and $K$ are the number of layers and constellation size, respectively. On the other hand, the SD exhibits ($\mathcal{O}(2^{K})$) for worst-case complexity. The FLOPs count for QPSK, 8PSK and 16 QAM constellation with varying transmitted symbol size, show that DetNet has lower FLOP count than those for semidefinite relaxation (SDR) based MIMO detection but has higher FLOP count than the AMP detector. However, the proposed network is not flexible to changes in the size of the constellation or the number of users. To address MIMO detection under imperfect CSI, He et al. introduced model-driven DL networks, orthogonal approximate message passing (OAMP) network, OAMP-Net [71], and its enhanced variant OAMP-Net2 [73]. OAMP-Net2 outperforms OAMP-Net by adapting to varying channel conditions. However, the robustness of the OAMP-Net2 under hardware impairment has not been verified in [73]. The OAMP algorithm is unfolded, and the linear MMSE (LMMSE) channel estimator is used to design the receiver in [73]. Learnable parameters, such as the step-size for updating the prior mean and variance in the MMSE estimator, are trained using SGD. These parameters in OAMP-Net2 help compensate for channel estimation errors, improving performance compared to DetNet [163].
@@
Furthermore, for overloaded massive MIMO systems, low-complexity detectors are essential. The trainable PGD (TPGD) algorithm [183] unfolds the PGD method with trainable parameters (step size) optimized via backpropagation and SGD. To further improve PGD with an increase in convergence speed, Yun et al. in [226] added momentum components to the iterative hidden layers. Momentum helps guide parameter updates in the right direction and reduces oscillations during PGD optimization [159]. Both networks in [183, 226] achieve lower complexity and outperform DetNet [163] and OAMP-Net2 [71]. A detector based on the Hubbard–Stratonovich transformation [182] avoids matrix inversion, offering faster training and lower execution cost compared to TPGD [183] and OAMP-Net2 [73]. MMNet [101] unfolds ISTA, alternating linear detection and nonlinear denoising, designed for ill-conditioned and non i.i.d. channels. MMNet balances complexity and flexibility, outperforming DetNet [163], OAMP-Net [71]. The MMNet outperforms the classical LMMSE detector with a performance gain of approximately 464-6 dB at a symbol error rate (SER) of $10^{-3}$ for different channel conditions.
@@
The CONCRETE MAP method in [17] solves the maximum a posteriori (MAP) detection problem via CONtinuous relaxation of disCRETE random variables, approximating MAP solutions using gradient descent unfolding. It generates soft outputs for integration with soft-input decoders and outperforms DetNet [163], OAMP-Net2 [73], and MMNet [101] in BER under correlated channels. In [204], conjugate gradient descent (CGD) iterations are unfolded to design the network, LcgNet. The authors also proposed a quantized version, QLcgNet, which minimizes memory usage with negligible performance degradation. LcgNet surpasses DetNet [163] in SER, complexity, and memory cost, making it suitable for massive MIMO system. Olutayo et al. [154] improved LcgNet by adding a preconditioner during training, enhancing convergence speed by optimizing the receiver filter’s spectrum. Preconditioning improves the condition number of a matrix [30], optimizing the receiver filter’s spectrum. This enhancement significantly increases the convergence speed compared to the original LcgNet in [204]. A dynamic conjugate gradient network further adapts to time-varying channels using self-supervised learning with forward error correction codes.
@@
Physical layer signal processing handles continuous-time analog waveforms, which must be quantized for hardware implementation. Low-resolution (LR) one-bit analog-to-digital converters (ADC) are widely used at radio frequency (RF) ends for high sampling rates, low cost, and low power [7]. Khobahi Khobahi et al. [102] developed an LR detection network (LoRDNet), which unfolds the PGD method for symbol detection using the ML estimator, learning channel behavior implicitly from training data without requiring CSI. LoRDNet outperforms classical ML [39] and DL networks like DeepSIC [175], even with small datasets ($\sim 500$ samples). A generalized PGD algorithm was proposed in [77], where multiple gradient steps are performed before projection to balance convergence and performance. The self-correcting auto detector (SAD) unfolds the PGD algorithm, by combining a self-correcting gradient module and a learnable denoising autoencoder projection. SAD [77] achieves better detection-complexity trade-off than DetNet [163], using only $4$ projection steps versus DetNet’s $15$ projection steps, significantly reducing latency and computational cost.
[…]
@@
TABLE III: Comparison of Classical vs Unfolded methods for Signal Detection.
$T$: Layers, $I$: Iterations, $M$: Symbol size, $N_{t}$: No. of transmit antenna, $N_{r}$: No. of receive antenna.
Unfolded Computational Complexity Number of Parameters (Unfolded) Convergence speed (Number of layers)
Model Classical Unfolded Simulation Setup Classical Unfolded
DetNet [163] Sphere Decoder: $\mathcal{O}(M^{N_{t}})$ $\mathcal{O}(TN_{t}^{2})$ $(6N_{r}N_{t}+2N_{r}+N_{r})T$ QPSK, 8-PSK, 16-QAM, $0.55$-Toeplitz channel, $N_{t}=\{15,20,30\}$, $N_{r}=\{20,25,30\}$. $\sim 1000$ iterations $30$ Layers
LcgNet [204] CGD: $\mathcal{O}(I(8N_{t}^{2}+14N_{t}+8))$ $\mathcal{O}(T(4N_{t}^{2}+6N_{t}))$ $(6N_{t}+4N_{t}^{2}+4T)$ MIMO $N_{r}=32$, $N_{t}=64$, Rayleigh fading channel. $\sim 40$ iterations 77 Layers
OAMP-Net2 [73] OAMP: $\mathcal{O}(N_{t}^{3})$ $\mathcal{O}(TN_{t}^{3})$ $4T$ MIMO Rayleigh fading channel, QPSK and 16-QAM, $N_{t}=N_{r}=4$ and $N_{t}=N_{r}=8$. $1000$ iterations 44 Layers
MMNet [101] MMSE: $\mathcal{O}(N_{r}^{3})$ $\mathcal{O}(N_{r}^{3})$ $2N_{t}(N_{r}+1)$ Massive MIMO, QAM modulation, 3GPP 3D channel, $N_{t}=\{16,32\}$, $N_{r}=64$. $50$ iterations $10$ Layers
ADMM-Net [103] ADMM: $\mathcal{O}(M^{3})$ $\mathcal{O}(M^{2})$ $5T$ MIMO Rayleigh fading channel, BPSK, 4-QAM and 16-QAM, $N_{t}=N_{r}=50$. $30$ iterations <20<20 Layers
@@
To avoid the computational burden incurred due to matrix inversion, the authors in [45] have proposed a sparse refinement architecture for symbol detection in uplink massive MIMO. The authors incorporated deep unfolding for components in matrix inversion, to reduce the detection complexity. The sparse weight matrix associated with the Jacobian iteration is trained across the layers of the network. Liao et al. [114] incorporated interference cancellation into DNN layers. Inspired by [137], their sequential detector trains residual error coefficients with MSE loss, reducing complexity by avoiding matrix inversion and relying on matrix addition and multiplication. Compared to DetNet [163], which requires $90$ layers for convergence, their model achieves convergence with just $8$ layers, offering significant computational simplicity. The authors in [103] developed an unfolded ADMM-based network for MIMO detection. They employed an inexact ADMM update as in [24, 155] to reduce complexity while achieving near-optimal performance. Compared to DetNet [163], the proposed network [103] requires fewer trainable parameters, resulting in improved performance in terms of symbol error. Since, in [103], perfect CSI was assumed for MIMO detection, the authors in [181], proposed an ADMM-based unfolded algorithm with channel estimation errors. A set of penalty and relaxation parameters are learned layer wise. Moreover, the complexity of the proposed network is reduced as compared to OAMP-Net2 [73]. The authors in [9] have proposed an unfolded coherent detector for optical communication. They have developed a MIMO filter architecture to compensate for hardware impairments such as transmitter and receiver in-phase/quadrature skew. The learnable parameters include the filter coefficients of the MIMO filter and are trained using the method in [183]. The results showed that the unfolded network performed well in the presence of chromatic dispersion, polarization rotation, and frequency offset accumulated over the fiber channel.
@@ III-B Modern Multicarrier Modulation
A cellular network must simultaneously support a growing number of active subscribers within a limited time-frequency spectrum [96]. By increasing bandwidth the symbol period is reduced which worsens the ISI in multi-path fading channels. OFDM mitigates ISI by splitting wideband signals into orthogonal narrowband subcarriers [208]. MIMO-OFDM detectors faces challenges due to high-dimensional signal space. An unfolded OAMP-based low-complexity receiver for cyclic prefix-free MIMO-OFDM is explored in [234]. The network was tested in over the air experiments and it performed reliably with low synchronization errors under imperfect CSI. Similarly, detection of coarsely quantized or LR signals in one-bit OFDM-based MIMO systems is a challenge. One-bit ADCs severely quantize received signals, discarding amplitude information, and hence, ML detection becomes intractable due to nonlinear quantization effects. Shao et al. [168] proposed an unfolded expectation maximization (EM) algorithm that iteratively estimates unquantized symbols, outperforming ZF detection at high SNR. Furthermore, Ullah et al. [191] unfolded a likelihood ascent search algorithm for joint detection and soft-output estimation. The likelihood ascent search algorithm is a neighborhood search-based detector for the MIMO system. The algorithm begins with an initial symbol estimate from linear detectors (ZF or MMSE) and iteratively refines the solution by flipping bits in the detected symbols to minimize the likelihood cost function. Near-optimal detection is achieved with the unfolded network, $1.2$ dB gap with 16-QAM as compared to SD at BER of $10^{-5}$ at polynomial complexity. Zhang et al. [235] unfolded expectation propagation (EP) with trainable layer-specific damping factors, providing robust detection under varying channels and noise, validated in over-the-air tests. In underwater acoustic communication, where radio or light waves fail, Zhao et al. in [238] proposed UDNet, an unfolded MMSE equalizer that generalizes well with low complexity in doubly-selective channels. In terms of performance, UDNet achieved better generalization and maintained lower computational complexity than DetNet [163]. Li et al. [110] proposed EP-Net for OTFS detection, training step-size, damping, and posterior distribution parameters to improve detection performance.
In [225], an unfolded OAMP-based detector for OTFS is designed to handle channel estimation errors. This detector is trained using imperfect CSI and optimizes three trainable parameters: step-sizes for updating the linear estimate, nonlinear estimate, and noise variance. Numerical analysis demonstrates that at SNR of $30$ dB, the proposed detector achieves a BER of $2.4\times 10^{-4}$ while EP-Net [110] achieves BER of $1.2\times 10^{-3}$.